### v1

# Self-playing Adversarial Language Game Enhances LLM Reasoning

Abstract
We explore the self-play training procedure of large language models (LLMs) in a two-player adversarial language game called Adversarial Taboo. In this game, an attacker and a defender communicate with respect to a target word only visible to the attacker. The attacker aims to induce the defender to utter the target word unconsciously, while the defender tries to infer the target word from the attacker’s utterances. To win the game, both players should have sufficient knowledge about the target word and high-level reasoning ability to infer and express in this information-reserved conversation. Hence, we are curious about whether LLMs’ reasoning ability can be further enhanced by Self-Play in this Adversarial language Game (SPAG). With this goal, we let LLMs act as the attacker and play with a copy of itself as the defender on an extensive range of target words. Through reinforcement learning on the game outcomes, we observe that the LLMs’ performance uniformly improves on a broad range of reasoning benchmarks. Furthermore, iteratively adopting this self-play process can continuously promote LLM’s reasoning ability. The code is at https://github.com/Linear95/SPAG.
Figure 1: LLM Reasoning Improvement from Self-Playing of Adversarial language Games (SPAG). With the epoch of SPAG increasing, the LLM reasoning ability continuously improves. Each axis is normalized by the maximum value.
1 Introduction
Large language models (LLMs), such as GPT-4 (OpenAI, 2023b) and Gemini (Team et al., 2023), have reformed the domain of artificial intelligence (AI) with astonishing language capacities, such as natural language understanding (Yang et al., 2023b; Touvron et al., 2023), text generation (Kocoń et al., 2023; Anil et al., 2023), machine translation (Jiao et al., 2023), and programming (Surameery and Shakor, 2023; Tian et al., 2023). However, the reasoning ability of LLMs, which is essential for complex problem-solving (Pan et al., 2023) and advanced intelligence-developing (Yao et al., 2021), still retains being challenged by various criteria including correctness (Zhang et al., 2023) and faithfulness (Turpin et al., 2024).
To address the reasoning challenge in LLMs, plenty of works have contributed in-depth efforts from the perspectives of Chain-of-Thought (CoT) prompt engineering(Wei et al., 2022; Yao et al., 2024; Ding et al., 2023), and the usage of auxiliary reasoning tools (Pan et al., 2023). However, both prompt-based and tool-calling methods require additional prompt designs, which are inconsistent and sensitive to different prompt patterns and LLM checkpoints (Turpin et al., 2024; Chu et al., 2023). More fundamental and consistent reasoning-improving approaches are based on post-pretraining (Azerbayev et al., 2023) and fine-tuning (Dong et al., 2023a). Nevertheless, these methods demand adequate high-quality textual data, which are inevitable to the massive costs of human annotation efforts (Singh et al., 2023).
To improve the reasoning ability of LLMs more efficiently, self-improvement methods, which enhance LLMs with model-generated synthetic data, have recently attracted increasing research attention (Singh et al., 2023; Huang et al., 2023; Burns et al., 2023; Cheng et al., 2023b; Chen et al., 2024). Self-improvement methods usually utilize the intrinsic language capability of LLMs to judge (Huang et al., 2023), filter (Yuan et al., 2024), or revise (Yuan et al., 2024) self-generated samples to enhance their quality. However, most self-improvement methods rely on a broad range of high-quality question queries to prevent over-fitting into a sub-domain of reasoning tasks, which still requires additional data collection and cleaning endeavors. Besides, the judgments from LLMs are not guaranteed objective. If an LLM already has an incorrect or biased recognition of a particular concept, the self-improvement process can reinforce and amplify the LLM’s cognitive dissonance.
Towards more general and objective self-reasoning-improvement methods, we are inspired by the advancement from AlphaGO (Silver et al., 2016) to AlphaGO Zero (Silver et al., 2017), in which the player’s intelligence continuously promotes without any human knowledge via self-play. Analogically, we expect to set up a language game where LLMs can improve their reasoning capacities from self-play. Although language games have attracted increasing attention (Lewis et al., 2017; Hausknecht et al., 2020; Abdulhai et al., 2023; Ma et al., 2023; Xu et al., 2023; Wu et al., 2024), most of them are specially designed with many customized rules, in lack of the generalization ability to improve the general capacities of LLMs. Among a few general-target language games including red-teaming (Ma et al., 2023), negotiation (Lewis et al., 2017), and bargain (Abdulhai et al., 2023), additional human judgments or reward models are required for outcome determination, which posts challenges on the efficiency and effectiveness of large-scale self-play sampling.
Based on the above consideration, we choose an adversarial language game called Adversarial Taboo (Yao et al., 2021), in which an attacker and a defender perform a conversation around a target word only visible to the attacker. The attacker aims to induce the defender to speak out the target word unconsciously; the defender tries to avoid unconscious utterance of the target word and guess the word from the dialogue history. By collecting target words from a vast vocabulary, this game can cover a broad range of topics providing sufficient generality for self-play. Besides, the game outcomes can be automatically and explicitly judged: we only need to check whether the target word appears in the defender’s utterances or inference patterns. We conduct the self-play of this adversarial language game using open-sourced LLMs, LLaMA-2-7B (Touvron et al., 2023) and Baichuan-2-13B (Yang et al., 2023a), with the target word selected from the list of 50K top-frequent words (Davies, 2020). Next, we conduct offline reinforcement learning (RL) on the game outcomes and observe significant performance improvement on a broad range of LLM reasoning benchmarks. Moreover, we iterate this sampling-learning process with three epochs, within which the LLMs’ reasoning can continuously obtain improvement. We believe this novel LLM training scheme, Self-Play from Adversarial language Game (SPAG), has great potential for developing advanced LLM capacities.
2 Preliminary
With the development of LLMs (OpenAI, 2023a, b), reinforcement learning (RL) has played an increasingly important role in language model training (Ouyang et al., 2022; Ramamurthy et al., 2023). The prime application scenario for RL in LLM training is reinforcement learning from human feedback (RLHF) (Yuan et al., 2023; Dong et al., 2023b; Cheng et al., 2023b). RLHF first learns a reward model $r({\bm{x}},{\bm{y}})$ from the human feedback preference pairs (Cheng et al., 2023a), and then optimizes the LLM policy $\pi_{\theta}({\bm{y}}|{\bm{x}})$ to maximize the expected reward value (Dong et al., 2023b):
${\mathcal{L}}_{\text{RLHF}}(\pi_{\theta})=-{\mathbb{E}}_{{\bm{x}}\sim{\mathcal{D}},{\bm{y}}\sim\pi_{\theta}({\bm{y}}|{\bm{x}})}[r({\bm{x}},{\bm{y}})],$ (1)
where $\text{KL}[\pi_{\theta}({\bm{y}}|{\bm{x}})\|\pi_{\text{ref}}({\bm{y}}|{\bm{x}})]$ is the Kullback-Leibler (KL) (Kullback, 1997) divergence between LLM policy $\pi_{\theta}({\bm{y}}|{\bm{x}})$ with a reference $\pi_{\text{ref}}({\bm{y}}|{\bm{x}})$. To learn the above objective, proximal policy optimization (PPO) (Schulman et al., 2017) algorithm has been recognized as the mainstream solution. In each update to equation 1, PPO minimizes:
$-{\mathbb{E}}_{{\bm{x}}\sim{\mathcal{D}},{\bm{y}}\sim\pi_{\bar{\theta}}({\bm{y}}|{\bm{x}})}\left[\frac{\pi_{\theta}({\bm{y}}|{\bm{x}})}{\pi_{\bar{\theta}}({\bm{y}}|{\bm{x}})}\hat{A}^{\pi_{\bar{\theta}}}-\beta\text{KL}[\pi_{\bar{\theta}}\|\pi_{\theta}]\right],$ (2)
where $\pi_{\bar{\theta}}$ is a copy of $\pi_{\theta}$ before the update, $\hat{A}^{\pi_{\bar{\theta}}}$ is the estimated advantage value (Schulman et al., 2016) with respect to the sampling policy $\pi_{\bar{\theta}}$. However, PPO for LLMs has been continually challenged due to its inefficient natural-text online sampling and the unstable training processes (Baheti et al., 2024). Among the improvements to PPO (Rafailov et al., 2023; Yuan et al., 2023; Dong et al., 2023b), Baheti et al. (2024) adopts the PPO objective into an offline scheme by using importance sampling (Neal, 2001), which is named Advantage-Leftover-Lunch (A-LoL):
$\nabla_{\theta}\hat{{\mathcal{L}}}_{\text{RLHF-A-LoL}}={\mathbb{E}}_{{\bm{x}}\sim{\mathcal{D}},{\bm{y}}\sim\pi_{\text{ref}}({\bm{y}}|{\bm{x}})}\left[\hat{A}^{\pi_{\text{ref}}}\cdot\frac{\pi_{\theta}({\bm{y}}|{\bm{x}})}{\pi_{\text{ref}}({\bm{y}}|{\bm{x}})}\cdot\nabla_{\theta}\log\pi_{\theta}({\bm{y}}|{\bm{x}})\right].$ (3)
Here the sample ${\bm{y}}\sim\pi_{\text{ref}}({\bm{y}}|{\bm{x}})$ and advantage $\hat{A}^{\pi_{\text{ref}}}$ are both from the reference distribution $\pi_{\text{ref}}({\bm{y}}|{\bm{x}})$ and calculated offline. Besides, Gulcehre et al. (2023) proposed Reinforced Self-Training (ReST) to simplify the RLHF scheme. With a threshold $\tau\in\mathbb{R}$, ReST updates the LLM by the reinforcement on the selected samples ${\mathcal{D}}_{\tau}=\{({\bm{x}},{\bm{y}}):r({\bm{x}},{\bm{y}})>\tau\}$:
${\mathcal{L}}_{\text{ReST}}(\pi_{\theta})={\mathbb{E}}_{{\bm{x}}\sim{\mathcal{D}},{\bm{y}}\sim\pi_{\text{ref}}({\bm{y}}|{\bm{x}})}\left[\bm{1}_{r({\bm{x}},{\bm{y}})>\tau}\cdot{\mathcal{L}}_{\theta}({\bm{x}},{\bm{y}})\right]={\mathbb{E}}_{{\mathcal{D}}_{\tau}}[{\mathcal{L}}_{\theta}({\bm{x}},{\bm{y}})],$ (4)
where ${\mathcal{L}}_{\theta}({\bm{x}},{\bm{y}})$ could be any offline RL loss such as A-LoL or the naive language modeling loss.
Figure 2: Examples of Adversarial Taboo with the same target word “conversation”. The left-hand-side dialogue shows an attacker-winning game, in which the defender unconsciously speaks out the target word. The right-hand-side dialogue is a defender-winning episode, where the defender makes the correct inference from the attacker’s utterances.
3 Self-play of Adversarial Language Games
The game of Adversarial Taboo is first introduced by Yao et al. (2021), in which an attacker $\mu$ and a defender $\nu$ involve an adversarial conversation. At the beginning of the game, the attacker is assigned a target word $w\in{\mathcal{V}}_{\text{target}}$, which is not informed to the defender. The attacker’s target is to induce the defender to speak the target word unconsciously. To achieve this goal, the attacker can talk about any topic related to $w$, except directly speaking out the target word. In contrast, the defender is required to infer the target word without any unconscious utterance of the word. If the defender has achieved sufficient confidence to infer the word, he/she can yell “I know the word! It is {target word} !”. Then the game terminates. If the guess is correct, the defender wins, otherwise the attacker wins. Besides, the game has a maximum number of turns $T_{0}$. If nobody wins during $T_{0}$ turns, there is a tie. The examples of Adversarial Taboo are shown in Figure 2.
3.1 Adversarial Language Game Modeling
We view the Adversarial Taboo as a two-player zero-sum Markov game (Littman, 1994), which can be described by a tuple as $({\mathcal{S}},{\mathcal{A}},F,r)$:
  • The state space ${\mathcal{S}}=\{{\bm{s}}_{t},{\bm{s}}^{\prime}_{t}|1\leq t\leq T_{0}\}$ contains two types of states, ${\bm{s}}^{\prime}_{t}=(w,{\bm{u}}_{1},{\bm{v}}_{1},{\bm{u}}_{2},\dots,{\bm{u}}_{t})$ and ${\bm{s}}_{t}=(w,{\bm{u}}_{1},{\bm{v}}_{1},{\bm{u}}_{2},\dots,{\bm{u}}_{t},{\bm{v}}_{t})$, where $\{{\bm{u}}_{i}\}_{i=1}^{t}$ and $\{{\bm{v}}_{i}\}_{i=1}^{t}$ are the utterances of the attacker and defender, respectively. Games start at ${\bm{s}}_{0}=(w)$ with a target word $w\in{\mathcal{V}}$ and end with $T_{0}$ maximum turns. States ${\bm{s}}^{\prime}_{t}$ and ${\bm{s}}_{t}$ end with utterances ${\bm{u}}_{t}$ and ${\bm{v}}_{t}$ for the defender and attacker to act, respectively.
  • The action space ${\mathcal{A}}$ is shared with both the attacker and defender, which is equivalent to the token sequence space of natural language ${\mathcal{N}}=\{{\bm{x}}=(x_{1},x_{2},\dots,x_{L})|x_{l}\in{\mathcal{V}},L\in\mathbb{N}_{+}\}$.
  • The transition function $F:{\mathcal{S}}\times{\mathcal{A}}\rightarrow{\mathcal{S}}$ deterministically appends the utterance ${\bm{u}}_{i}$ or ${\bm{v}}_{i}$ at the end of the dialogue, and converts ${\bm{s}}^{\prime}_{t}=F({\bm{s}}_{t-1},{\bm{u}}_{t})$ and ${\bm{s}}_{t}=F({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})$.
  • The reward $r:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\mathbb{R}$ evaluates the actions ${\bm{u}}_{i},{\bm{v}}_{i}\in{\mathcal{A}}$ based on their corresponding states ${\bm{s}},{\bm{s}}^{\prime}\in{\mathcal{S}}$ with rewards $r({\bm{s}},{\bm{u}})$ and $r({\bm{s}}^{\prime},{\bm{v}})$, respectively. Given a game episode $\bm{\tau}=({\bm{s}}_{0},{\bm{s}}^{\prime}_{1},{\bm{s}}_{1},\dots,{\bm{s}}^{\prime}_{T},{\bm{s}}_{T})$, we denote the attacker total reward $R(\bm{\tau})=\sum_{t=1}^{T}r({\bm{s}}_{t-1},{\bm{u}}_{t})$, so the defender total reward $\sum_{t=1}^{T}r({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})=-R(\bm{\tau})$ in this zero-sum game. More detailed reward designs with heuristic rules for the Adversarial Taboo can be found in Appendix B.
In the above language game, we denote $\mu({\bm{u}}|{\bm{s}})$ and $\nu({\bm{v}}|{\bm{s}}^{\prime})$ as the attacker’s policy and the defender’s policy, respectively. Then each episode $\bm{\tau}$ can be regarded as a sample from the following distribution:
$(\mu\times\nu):=P(\bm{\tau})=P({\bm{s}}_{0})\prod_{t=1}^{T}P({\bm{s}}^{\prime}_{t}|{\bm{s}}_{t-1})\prod_{t=1}^{T}P({\bm{s}}_{t}|{\bm{s}}^{\prime}_{t})=P(w)\prod_{t=1}^{T}\mu({\bm{u}}_{t}|{\bm{s}}_{t-1})\prod_{t=1}^{T}\nu({\bm{v}}_{t}|{\bm{s}}^{\prime}_{t}).$ (5)
Hence, the objective of the adversarial language self-play is:
$\max_{\mu}\min_{\nu}{\mathcal{L}}_{\text{AG}}(\mu,\nu):=\mathbb{E}_{{\bm{\tau}}\sim{\mu\times\nu}}[R(\bm{\tau})].$ (6)
To play the above adversarial game with an LLM generation policy $\pi_{\theta}({\bm{y}}|{\bm{x}})$, we first designed prompt templates $f_{\text{attack}},f_{\text{defend}}:{\mathcal{S}}\rightarrow{\mathcal{N}}$ for the attacker and defender respectively to convert the game states into natural language task descriptions. Next, we introduce the game policies for the two players:
$\mu_{\theta}({\bm{u}}|{\bm{s}})=\pi_{\theta}({\bm{u}}|f_{\text{attack}}({\bm{s}})),\text{ and \ }\nu_{\theta}({\bm{v}}|{\bm{s}}^{\prime})=\pi_{\theta}({\bm{v}}|f_{\text{defend}}({\bm{s}}^{\prime})).$ (7)
The detailed prompt templates for the game are demonstrated in Appendix A.
  Inputs: LLM policy $\pi_{\theta}({\bm{y}}|{\bm{x}})$, target word $w$, attacker and defender prompts $f_{\text{attack}},f_{\text{defend}}$.
  Set the initial state ${\bm{s}}_{0}=(w)$.
  for $t$ from $1$ to $T$ do
     Sample an attacker utterance ${\bm{u}}_{t}\sim\mu_{\theta}({\bm{u}}_{t}|{\bm{s}}_{t-1})=\pi_{\theta}({\bm{y}}={\bm{u}}_{t}|{\bm{x}}=f_{\text{attack}}({\bm{s}}_{t-1}))$.
     Update state ${\bm{s}}^{\prime}_{t}=(w,{\bm{u}}_{1},{\bm{v}}_{1},\dots,{\bm{u}}_{t-1},{\bm{v}}_{t-1},{\bm{u}}_{t})$.
     Sample a defender utterance ${\bm{v}}_{t}\sim\nu_{\theta}({\bm{v}}_{t}|{\bm{s}}^{\prime}_{t})=\pi_{\theta}({\bm{y}}={\bm{v}}_{t}|{\bm{x}}=f_{\text{defend}}({\bm{s}}^{\prime}_{t}))$.
     Update state ${\bm{s}}_{t}=(w,{\bm{u}}_{1},{\bm{v}}_{1},\dots,{\bm{u}}_{t-1},{\bm{v}}_{t-1},{\bm{u}}_{t},{\bm{v}}_{t})$.
  end for
  Collect an episode $\tau=({\bm{s}}_{0},{\bm{s}}^{\prime}_{1},{\bm{s}}_{1},\dots,{\bm{s}}^{\prime}_{T},{\bm{s}}_{T})$
Algorithm 1 Data collection of LLM self-plays for the adversarial language game.
3.2 Imitation Learning
Due to the capability limitations of the open-source LLMs, the generation policy $\pi_{\theta}({\bm{y}}|{\bm{x}})$ can not guarantee to follow the instructions of the game prompts $f_{\text{attack}}({\bm{s}})$ and $f_{\text{defend}}({\bm{s}}^{\prime})$. Therefore, before the self-play learning stage, we first conduct an imitation learning (behavior cloning) process to ensure the LLM behaviors $\pi_{\theta}({\bm{u}}|f_{\text{attack}}({\bm{s}}))$ and $\pi_{\theta}({\bm{v}}|f_{\text{defend}}({\bm{s}}^{\prime}))$ consistent with game rules.
We collect a group of game trajectories ${\mathcal{T}}_{\text{im}}$ from the self-plays of GPT-4 (Achiam et al., 2023) and divide it into an attacker-winning set ${\mathcal{T}}^{\text{attack}}_{\text{im}}=\{\bm{\tau}\in{\mathcal{T}}_{\text{im}}:R(\bm{\tau})>0\}$ and a defender-winning set ${\mathcal{T}}^{\text{defend}}_{\text{im}}=\{\bm{\tau}\in{\mathcal{T}}_{\text{im}}:R(\bm{\tau})<0\}$. The imitation learning loss is to maximize the log-likelihood of winners’ actions:
${\mathcal{L}}_{\text{im}}^{\text{attack}}(\pi_{\theta})=$ $-\mathbb{E}_{\bm{\tau}\in{\mathcal{T}}^{\text{attack}}_{\text{im}}}\left[\frac{1}{T}\sum_{t=1}^{T}\log\pi_{\theta}({\bm{u}}_{t}|f_{\text{attack}}({\bm{s}}_{t-1}))+\beta_{1}\text{KL}[\pi_{\theta}\|\pi_{\text{ref}}]\right],$ (8)
${\mathcal{L}}_{\text{im}}^{\text{defend}}(\pi_{\theta})=$ $-\mathbb{E}_{\bm{\tau}\in{\mathcal{T}}^{\text{defend}}_{\text{im}}}\left[\frac{1}{T}\sum_{t=1}^{T}\log\pi_{\theta}({\bm{v}}_{t}|f_{\text{defend}}({\bm{s}}^{\prime}_{t}))+\beta_{1}\text{KL}[\pi_{\theta}\|\pi_{\text{ref}}]\right],$ (9)
where the regularizer $\text{KL}(\pi_{\theta}\|\pi_{\text{ref}})$ prevents the model from over-fitting on the language gaming task and decaying of the general language abilities. Then, the overall imitation learning objective is
${\mathcal{L}}_{\text{im}}(\pi_{\theta})=\frac{1}{2}\cdot{\mathcal{L}}_{\text{im}}^{\text{attack}}(\pi_{\theta})+\frac{1}{2}\cdot{\mathcal{L}}_{\text{im}}^{\text{defend}}(\pi_{\theta}).$ (10)
3.3 Reinforcement Learning from Self-play
Imitation learning enables LLM to behave correctly following the game rules, then we conduct the self-plays by letting the LLM $\pi_{\theta}({\bm{y}}|{\bm{x}})$ play as the attacker $\mu_{\theta}({\bm{u}}|{\bm{s}})=\pi_{\theta}({\bm{u}}|f_{\text{attack}}({\bm{s}}))$ and the defender $\nu_{\theta}({\bm{v}}|{\bm{s}}^{\prime})=\pi_{\theta}({\bm{v}}|f_{\text{defend}}({\bm{s}}^{\prime}))$ alternatively. Note that the self-play sampling process involves massive multi-turn auto-regressive text generation of the LLM, which causes heavy computational complexity and makes the online RL training intensively inefficient. Therefore, we use an offline learning scheme: (1) make a copy $\pi_{\bar{\theta}}$ of current LLM policy $\pi_{\theta}$; (2) collect self-play episodes ${\mathcal{T}}_{\bar{\theta}}=\{\bm{\tau}\sim\mu_{\bar{\theta}}\times\nu_{\bar{\theta}}\}$ from games between attacker $\mu_{\bar{\theta}}$ and defender $\nu_{\bar{\theta}}$; (3) update $\pi_{\theta}$ via reinforcement learning from ${\mathcal{T}}_{\bar{\theta}}$. The details of self-play ${\mathcal{T}}_{\bar{\theta}}$ collection are shown in Algorithm 1.
With a group of episodes ${\mathcal{T}}_{\bar{\theta}}$, we first consider the learning objective from the view of attacker $\mu_{\theta}$ with defender $\nu_{\bar{\theta}}$ fixed as ${\mathcal{L}}_{\text{AG}}(\mu_{\theta},\nu_{\bar{\theta}})={\mathbb{E}}_{\bm{\tau}\sim\mu_{\theta}\times\nu_{\bar{\theta}}}[R(\bm{\tau})]$. Then we calculate the corresponding policy gradient for the attacker as:
$\nabla_{\theta}{\mathcal{L}}_{\text{AG}}(\mu_{\theta},\nu_{\bar{\theta}})=$ ${\mathbb{E}}_{\bm{\tau}\sim\mu_{\theta}\times\nu_{\bar{\theta}}}\left[\sum_{t=1}^{T}A^{\mu_{\theta}}_{t}\cdot\nabla_{\theta}\log\mu_{\theta}({\bm{u}}_{t}|{\bm{s}}_{t-1})\right]$
$=$ ${\mathbb{E}}_{\bm{\tau}\sim\mu_{\bar{\theta}}\times\nu_{\bar{\theta}}}\left[\sum_{t=1}^{T}A^{\mu_{\bar{\theta}}}_{t}\cdot\frac{\mu_{\theta}({\bm{u}}_{t}|{\bm{s}}_{t-1})}{\mu_{\bar{\theta}}({\bm{u}}_{t}|{\bm{s}}_{t-1})}\cdot\nabla_{\theta}\log\mu_{\theta}({\bm{u}}_{t}|{\bm{s}}_{t-1})\right],$ (11)
where $A^{\mu_{\theta}}_{t}=A^{\mu_{\theta}}({\bm{s}}_{t-1},{\bm{u}}_{t})$ and $A^{\mu_{\bar{\theta}}}_{t}=A^{\mu_{\bar{\theta}}}({\bm{s}}_{t-1},{\bm{u}}_{t})$ are the advantage values for the attacker in the self-plays $\mu_{\theta}\times\nu_{\bar{\theta}}$ and $\mu_{\bar{\theta}}\times\nu_{\bar{\theta}}$, respectively. The detailed estimation of $A^{\mu_{\bar{\theta}}}_{t}$ are described in the Appendix D. In equation 3.3, we apply the importance sampling strategy to estimate the expectation of $\mu_{\theta}({\bm{u}}_{t}|{\bm{s}}_{t-1})$ with the sampled actions of $\mu_{\bar{\theta}}({\bm{u}}_{t}|{\bm{s}}_{t-1})$ in ${\mathcal{T}}_{\bar{\theta}}$. Inspired by TRPO (Schulman et al., 2015) and PPO (Schulman et al., 2017), we design the following learning loss to optimize the objective ${\mathcal{L}}_{\text{AG}}(\mu_{\theta},\nu_{\bar{\theta}})$ with a KL regularizer $\text{KL}(\pi_{\theta}\|\pi_{\bar{\theta}})$ to prevent the LLM from overfitting, :
${\mathcal{L}}^{\text{attack}}_{\text{sp}}(\pi_{\theta})=$ $-{\mathbb{E}}_{\bm{\tau}\in{\mathcal{T}}_{\bar{\theta}}}\left[\sum_{t=1}^{T}\frac{\mu_{\theta}({\bm{u}}_{t}|{\bm{s}}_{t-1})}{\mu_{\bar{\theta}}({\bm{u}}_{t}|{\bm{s}}_{t-1})}A^{\mu_{\bar{\theta}}}_{t}-\beta_{2}\text{KL}[\pi_{\theta}\|\pi_{\bar{\theta}}]\right]$
$=$ $-{\mathbb{E}}_{\bm{\tau}\in{\mathcal{T}}_{\bar{\theta}}}\left[\sum_{t=1}^{T}\frac{\pi_{\theta}({\bm{u}}_{t}|f_{\text{attack}}({\bm{s}}_{t-1}))}{\pi_{\bar{\theta}}({\bm{u}}_{t}|f_{\text{attack}}({\bm{s}}_{t-1}))}A^{\mu_{\bar{\theta}}}_{t}-\beta_{2}\text{KL}[\pi_{\theta}\|\pi_{\bar{\theta}}]\right].$ (12)
Similarly, from the perspective of the defender, we can design the corresponding loss:
${\mathcal{L}}^{\text{defend}}_{\text{sp}}(\pi_{\theta})=-{\mathbb{E}}_{\bm{\tau}\in{\mathcal{T}}_{\bar{\theta}}}\left[\sum_{t=1}^{T}\frac{\pi_{\theta}({\bm{v}}_{t}|f_{\text{defend}}({\bm{s}}^{\prime}_{t}))}{\pi_{\bar{\theta}}({\bm{v}}_{t}|f_{\text{defend}}({\bm{s}}^{\prime}_{t}))}A^{\nu_{\bar{\theta}}}_{t}-\beta_{2}\text{KL}[\pi_{\theta}\|\pi_{\bar{\theta}}]\right].$ (13)
In practice, the above policy gradient method can cause training instability with negative advantage values, because of the exposure bias of auto-regressive language model (Bengio et al., 2015): losses ${\mathcal{L}}^{\text{attack}}_{\text{sp}}(\pi_{\theta})$ and ${\mathcal{L}}^{\text{defend}}_{\text{sp}}(\pi_{\theta})$ treat each entire utterance as an action, while each utterance is auto-regressively generated by the LLM $\pi_{\theta}$ via the next-token prediction. With a negative advantage as the multiplier, the language modeling loss can damage the LLM’s language capacity instead of punishing the ineffective game strategies. To mitigate the issue, we utilize the methodology in ReST (Gulcehre et al., 2023), which considers the offline learning with samples $\{\bm{\tau}:R(\bm{\tau})>\xi\}$ selected by a reward threshold $\xi$. In the self-play setup, we set the reward threshold $\xi=0$ and select the attacker-winning episodes ${\mathcal{T}}^{\text{attack}}_{\bar{\theta}}=\{\bm{\tau}\in{\mathcal{T}}_{\bar{\theta}}:R(\bm{\tau})>0\}$ and the defender-winning episodes ${\mathcal{T}}^{\text{defend}}_{\bar{\theta}}=\{\bm{\tau}\in{\mathcal{T}}_{\bar{\theta}}:R(\bm{\tau})<0\}$ for the attacker and the defender training respectively. Hence the final self-play of adversarial language games (SPAG) objective is:
${\mathcal{L}}_{\text{SPAG}}(\pi_{\theta})=$ $-\frac{1}{2}{\mathbb{E}}_{{\mathcal{T}}_{\bar{\theta}}^{\text{attack}}}\left[\sum_{t=1}^{T}\frac{\pi_{\theta}({\bm{u}}_{t}|f_{\text{attack}}({\bm{s}}_{t-1}))}{\pi_{\bar{\theta}}({\bm{u}}_{t}|f_{\text{attack}}({\bm{s}}_{t-1}))}A^{\mu_{\bar{\theta}}}_{t}-\beta_{2}\text{KL}[\pi_{\theta}\|\pi_{\bar{\theta}}]\right]$ (14)
$-\frac{1}{2}$ ${\mathbb{E}}_{{\mathcal{T}}_{\bar{\theta}}^{\text{defend}}}\left[\sum_{t=1}^{T}\frac{\pi_{\theta}({\bm{v}}_{t}|f_{\text{defend}}({\bm{s}}^{\prime}_{t}))}{\pi_{\bar{\theta}}({\bm{v}}_{t}|f_{\text{defend}}({\bm{s}}^{\prime}_{t}))}A^{\nu_{\bar{\theta}}}_{t}-\beta_{2}\text{KL}[\pi_{\theta}\|\pi_{\bar{\theta}}]\right]-\alpha{\mathbb{E}}_{({\bm{x}},{\bm{y}})\sim{\mathcal{D}}_{\text{SFT}}}[\log\pi_{\theta}({\bm{y}}|{\bm{x}})],$
where ${\mathbb{E}}_{{\mathcal{D}}_{\text{SFT}}}[\log\pi_{\theta}({\bm{y}}|{\bm{x}})]$ is the log-likelihood on an SFT dataset ${\mathcal{D}}_{\text{SFT}}$ to prevent the SPAG optimization from losing general LM abilities.
  Inputs: LLM policy $\pi_{\theta}({\bm{y}}|{\bm{x}})$, target word set ${\mathcal{V}}_{\text{target}}$, attacker and defender prompts $f_{\text{attack}},f_{\text{defend}}$.
  for self-play epoch iterations do
     Make a copy of $\pi_{\theta}$ as $\pi_{\bar{\theta}}$.
     Set $\mu_{\bar{\theta}}({\bm{u}}|{\bm{s}})=\pi_{\bar{\theta}}({\bm{u}}|f_{\text{attack}}({\bm{s}}))$ and $\nu_{\bar{\theta}}({\bm{v}}|{\bm{s}}^{\prime})=\pi_{\bar{\theta}}({\bm{v}}|f_{\text{defend}}({\bm{s}}^{\prime}))$.
     For each $w\in{\mathcal{V}}_{\text{target}}$, sample a episode $\bm{\tau}\sim\mu_{\bar{\theta}}\times\nu_{\bar{\theta}}$. Collect ${\mathcal{T}}_{\bar{\theta}}=\{\bm{\tau}\sim\mu_{\bar{\theta}}\times\nu_{\bar{\theta}}\}$
     Split the attacker-winning set ${\mathcal{T}}_{\bar{\theta}}^{\text{attack}}$ and the defender-winning set ${\mathcal{T}}_{\bar{\theta}}^{\text{defend}}$.
     Update $\pi_{\theta}$ with loss ${\mathcal{L}}_{\text{SPAG}}(\pi_{\theta})$.
  end for
Algorithm 2 Self-play of adversarial language games (SPAG)
4 Experiments
To verify the effectiveness of SPAG, we select open-source pretrained LLMs of different sources and model sizes, particularly LLaMA-2-7B (Touvron et al., 2023) and Baichuan-2-13B (Yang et al., 2023a). The evaluation metrics include the commonly used reasoning benchmarks (Gao et al., 2023) and the game-play win rates between different self-play checkpoints and GPT-4 (OpenAI, 2023b).
4.1 Experimental Setups
Training Data Preparation
The training data preparation consists of the following three parts. More data collection details are in Appendix A.
  • Target Words: We aim to play the adversarial game with an extensive range of target words, so that diverse topics of natural language can be discussed during the self-play processes, which enhances LLMs’ performance with rich generalization ability. Therefore, we collect the 50K most frequently used words from the Corpus of Contemporary American (CoCA) (Davies, 2020) as the self-play target word set ${\mathcal{V}}_{\text{target}}$. The stop words defined in NLTK (Bird, 2006), which are commonly used words with insignificant practical meaning, are filtered out of the target word set.
  • Imitation Learning Data: To collect the imitation learning data for enabling the instruction-following ability of open-source LLMs, we use the same self-play algorithm 1 via the GPT-4 (OpenAI, 2023b) API and play Adversarial Taboo with one episode per word. The attacker and defender prompts are in Appendix A.1. Due to the resource limitation, we only collect the GPT-4 self-plays with the top 30K words from ${\mathcal{V}}_{\text{target}}$. The results are collected as ${\mathcal{T}}_{\text{im}}$ for equation 10.
  • Supervised Funetuning (SFT) Data: We also prepare general query-response SFT data to avoid LLMs from being over-fitted on the adversarial game. We use Alpaca (Taori et al., 2023) as the SFT set, which contains 52K instruction-following data from GPT-3 (Brown et al., 2020).
Evaluation
To evaluate the LLMs’ performance, we consider the following two perspectives:
  • Reasoning Benchmarks: To test the reasoning ability, we consider the following commonly used benchmarks including BIG-Bench Hard (BBH) (Suzgun et al., 2022), ARC easy (ARC-e) & challenge (ARC-c) (Clark et al., 2018), Mutual (Cui et al., 2020), WinoGrande (Sakaguchi et al., 2019), LogiQA2 (Liu et al., 2023), PIQA (Bisk et al., 2020). BBH requires the exact match of the generated answers, the other benchmarks are all within the multiple-choice form. Besides reasoning metrics, we test MMLU (Hendrycks et al., 2020) as a general performance evaluation for LLMs. Additionally, we calculate the geometric mean of the numerical results of all benchmarks as an overall performance measurement. More benchmark details can be found in Appendix C.
  • Game Win Rates: Besides the general LLM capabilities, we report the win rates on a testing vocabulary ${\mathcal{V}}_{\text{test}}$ to validate if the game skills of LLMs improve through the reinforcement learning of self-play. ${\mathcal{V}}_{\text{test}}$ contains 150 typical words manually selected for different topics maintaining a strong diversity. We denote the number of winning episodes as $N_{\text{win}}$, the number of losing episodes as $N_{\text{lose}}$, and the number of tied games as $N_{\text{tie}}$. Then the player win rate is calculated as $(N_{\text{win}}+0.5N_{\text{tie}})/(N_{\text{win}}+N_{\text{loss }}+N_{\text{tie}})$. The invalid game episodes, where LLM players do not strictly follow the game rules, are ignored.
Training Details
Most of our training setups follow Alpaca (Taori et al., 2023). For imitation learning, the learning rate is 5e-6, and the KL-penalty coefficient $\beta_{1}=0.1$. For SPAG training, the learning rate is 2e-6, the KL-penalty coefficient $\beta_{2}=0.2$, and the SFT coefficient $\alpha=0.5$. For the Alpaca SFT baseline, we exactly follow the training setups of Alpaca and set the learning rate to 2e-6. Among all training stages, the batch size is 128 and the max sequence length is 2048. Each training process maintains one epoch. The reward decay weight $\gamma$ to the number of interaction turns is 0.8.
Table 1: Reasoning Performance of SPAG on LLaMA-2-7B.
MMLU BBH Mutual ARC-e ARC-c LGQA2 WGrande PIQA GEO-Mean
Base 45.80 32.48 50.90 76.30 43.26 25.32 69.14 78.07 49.17
AlpacaSFT-1 35.17 30.24 53.95 76.81 44.97 28.94 69.61 78.07 48.61
AlpacaSFT-2 44.17 32.50 55.08 77.15 46.50 29.20 68.67 78.24 50.82
AlpacaSFT-3 45.87 31.52 54.18 75.25 45.05 29.07 66.85 76.71 50.08
Imitation 46.15 32.74 52.82 76.81 44.80 27.10 69.46 78.24 50.22
SPAG-1 47.01 34.39 54.85 77.69 45.65 29.83 68.90 78.89 51.69
SPAG-2 47.28 34.73 54.97 78.45 46.84 30.08 69.61 79.33 52.19
SPAG-3 47.11 34.94 55.30 78.54 47.53 30.98 69.93 79.38 52.58
Table 2: Reasoning Performance of SPAG on Baichuan-2-13B.
MMLU BBH Mutual ARC-e ARC-c LGQA2 WGrande PIQA GEO-Mean
Base 59.00 39.03 53.72 77.36 46.84 30.73 69.93 77.42 54.21
AlpacaSFT-1 52.94 36.52 58.35 74.12 44.88 33.33 71.35 77.20 53.68
AlpacaSFT-2 51.27 36.60 57.67 73.36 44.28 33.46 69.77 75.90 53.00
AlpacaSFT-3 52.14 36.57 55.08 69.44 42.15 33.91 66.61 74.59 51.79
Imitation 58.37 39.49 57.11 76.60 47.53 33.65 70.59 78.49 55.45
SPAG-1 57.93 39.81 57.45 78.20 48.55 35.05 70.67 78.69 56.09
SPAG-2 57.99 39.97 57.67 78.32 49.83 35.62 71.03 78.83 56.52
SPAG-3 57.75 40.30 57.79 78.11 50.00 36.26 71.43 79.05 56.75
4.2 Result Analysis
The main results of our experiments are shown in Figure 1, where each axis is normalized by the maximum performance value, representing the reasoning performance on a particular benchmark. For notation simplification, we call the LLM after imitating learning as IM model, and the LLM after the $i$-th epoch of SPAG training as SPAG-$i$ ($i=1,2,3$).
By imitation learning on the collected GPT-4 self-play episodes, both LLaMA-2-7B and Baichuan-2-13B have obtained uniform performance improvements on all the reasoning benchmarks. As in Figure 1, both two gray regions are completely wrapped by the blue-line polygons. Besides, as for the general language capacity, the imitation-learned LLaMA-2 model achieves a better MMLU performance than the original LLaMA-2 base in Table 2. Although the imitation result of Baichuan-2-13B on MMLU slightly under-performs the base model, the numerical difference is insignificant compared to the reasoning gains, which means the imitation learning of the GPT-4 game behaviors can improve LLM reasoning while preserving the general language capacity.
After the imitation learning, we conduct three epochs of SPAG as described in Algorithm 2. As shown in Figure 1, on most reasoning benchmarks, both LLaMA-2-7B and Baichuan-2-13B have their performance steadily improved with the SPAG epoch increasing. For LLaMA-2-7B, although the first-epoch SPAG has a relatively lower performance than the behavior cloning model on the W, after an additional epoch of self-play iteration, SPAG-2 has achieved sufficient improvement to surpass the performance of the IM model. Considering the general LLM understanding performance on MMLU, SPAG models can not guarantee continuous improvement, especially for Baichuan-2-13B whose performance slightly decays during the SPAG training. Compared to the improvements in the reasoning benchmarks, this language-understanding decline is still within an acceptable range, for the overall performance (GEO-Mean) maintains increase significantly.
Since the SPAG training loss includes the SFT data, we conduct an ablation study to test whether the performance gains on reasoning benchmarks come from the SFT data or our SPAG method. More specifically, we follow the training setups in Alpaca (Taori et al., 2023) and conduct the supervised fine-tuning on both LLaMA-2-7B and Baichuan-2-13B with three epochs. The checkpoint after the $i$th-epoch training is denoted as AlpacaSFT-$i$ ($i=1,2,3$) and tested on the same evaluation sets. The SFT models’ performances are also reported in Figure 1 and Table 1 & 2. With the LLaMA base, our SPAG-3 model can uniformly defeat the SFT baselines with clear performance gaps on all benchmarks. With the Baichuan base, except the Mutual set, the SPAG models maintain noticeable advantages on other metrics. Considering the distinct surpass of overall performance, we claim that the major contribution to the reasoning improvements belongs to our self-play & reinforcement learning scheme.
Besides the natural language evaluation, we review the LLM performance as game players in terms of the win rates. We first test our IM models and SPAG models with GPT-4 as the opponent. For each testing word in ${\mathcal{V}}_{\text{test}}$, we play the language game twice, once GPT-4 as the attacker, once GPT-4 as the defender. The average win rates are reported in the left-hand-side plot of Figure 3, in which one can observe uniform and continuous win rate improvement of SPAG models playing with GPT-4. We also let SPAG models play the game against each other and report the attacker win rate in the right-hand-side plot of Figure 3, in which we find the defender’s performance significantly enhanced through the SPAG training. Self-play examples of different SPAG models are in Appendix E.
Figure 3: Left: average win rates of SPAG models playing against GPT-4. Right: average win rate of SPAG models playing as the attacker against different-epoch checkpoints.
5 Conclusion
To improve the reasoning capacity of LLMs more efficiently, we explore a novel training strategy called Self-Play learning in Adversarial language Game (SPAG). In our method, a given LLM first learns to act as an attacker and a defender to play the language game named Adversarial Taboo via imitation learning. Next, we collect self-play episodes from the LLM playing against a copy of itself in the adversarial game. Finally, the LLM is further reinforced on the selected winning episodes via our SPAG algorithm. We repeat this self-play & reinforcement process for three epochs and find that the LLM’s reasoning performance continuously and uniformly improves on various benchmarks. Our SPAG algorithm opens up a new path to improve the fundamental capabilities of LLMs from the perspective of multi-agent self-play. With more elaborate language game designs, we believe the self-play approaches have great potential for developing a broader range of advanced language abilities of LLMs.
Appendix A Data Collection Details
A.1 GPT-4 Self-play Data Collection
To collect GPT-4 self-play data for imitation learning, we use the following system prompts to let GPT-4 play as the attacker and defender for the next-turn utterances.
- Attacker System Prompt: {mdframed} Lets play a game. Do not mention you are an AI developed by OpenAI.
Descriptions of the game
          Game Name: Adversarial Taboo
          Rules:
                   1. There are two players in the game, one is called attacker, another is called defender.
                   2. At the begining, attacker will get a target word, it is talking with a defender who is not
                   informed with the target word. The task of the attacker is to induce the defender to utter
                   the target word unconciously, then the attacker win the game. At the same time, the
                   defender is guessing the target word, if the defender indentifies the target word, the
                   defender is win.
                   3. The game will totally process {max_turns} turns. if nobody wins during {max_turns}
                   turns, there is a tie.
Game states
          Role: You are attacker. Make sure you understand the rules of the game.
          Attention:
                   1. The defender only have one chance to guess the target word, so do not let him guess too
                   many times.
                   2. Defender must guess the word and its form correctly to win the game.
                   3. Do not utter the target word.
          Target word: {target_word}
- Defender System Prompt: {mdframed} Lets play a game. Do not mention you are an AI developed by OpenAI.
Descriptions of the game
          Game Name: Adversarial Taboo
          Rules:
                   1. There are two players in the game, one is called attacker, another is called defender.
                   2. At the begining, attacker will get a target word, it is talking with a defender who is not
                   informed with the target word. The task of the attacker is to induce the defender to utter
                   the target word unconciously, then the attacker win the game. At the same time, the
                   defender is guessing the target word, if the defender indentifies the target word, the
                   defender is win.
                   3. The game will totally process {max_turns} turns. if nobody wins during {max_turns}
                   turns, there is a tie.
Game states
          Role: Your are a defender. Make sure you understand the rules of the game.
          Attention:
                   1. Do not utter your thinking process.
                   2. You can only guess once.
                   3. If you have identified the target word, utter ”I know word! It is <target word>”,
                   rigorously in this form. If your target word is correct, you win the game, otherwise you
                   lose. You only have one chance, so be careful to utter your answer.
There are two slots in the system prompts: “max_turns” and “target_word”. “max_turns” is an integer number randomly selected from $[3,8]$. “target_word” is filled by the target word of the game.
A.2 Target Words Collection
We downloaded 50k words and corresponding word frequencies from the Corpus of Contemporary American [Davies, 2020]. Then, we sorted the words by frequency and removed the words with less than two letters. Next, we removed the stop words using the NLTK package. Finally, we selected the first 30k words from the sorted list. For the test words, we manually selected 150 words from diverse subjects, including economics, physics, mathematics, and medicine.
A.3 Training Prompt
For training and inference of IM models and SPAG models, we use the following attacker and defender prompts.
- Attacker Prompt: {mdframed} Play the game of Adversarial Taboo. In this game, there are two players, an attacker and a defender.
At the beginning, the attacker is assigned a target word, with which the defender is not informed. The task of the attacker is to induce the defender to utter the target word unconsciously, then the attacker wins the game. However, the attacker is not allowed to utter the target word, otherwise the attacker loses.
At the same time, the defender tries to figure out the target word. If the defender identifies the target word, he can say ”I know the word! It is ‘target word‘!”. Then the defender wins. Remember, the defender can only guess the word once, otherwise, he will directly lose the game.
The game has {max_turns} turns. if nobody wins during {max_turns} turns, there is a tie.
### Game History: {history}
### Instruction: You are the attacker. The target word is ‘{target_word}‘. Provide your response for the next turn.
### Response:
- Defender Prompt: {mdframed} Play the game of Adversarial Taboo. In this game, there are two players, an attacker and a defender.
At the beginning, the attacker is assigned a target word, which the defender is not informed of. The task of the attacker is to induce the defender to utter the target word unconsciously, then the attacker wins the game. However, the attacker is not allowed to utter the target word, otherwise, the attacker loses.
At the same time, the defender tries to figure out the target word. If the defender identifies the target word, he can say ”I know the word! It is ‘target word‘!”. Then the defender wins. Remember, the defender can only guess the word once, otherwise, he will directly lose the game.
The game has {max_turns} turns. if nobody wins during {max_turns} turns, there is a tie.
### Game History: {history}
### Instruction: You are the defender. Provide your response to infer the target word.
### Response:
To prevent LLMs from being over-fitted to a single prompt template during the self-play training, we rewrite the above prompt into 8 different expressions via GPT-4. Besides, we randomly select the instruction-tuning and chat forms to make up the SFT data.
Appendix B Reward Design Details
We heuristically design rewards to ensure that $R(\bm{\tau})=1$ if the attacker wins, $R(\bm{\tau})=-1$ if the defender wins, and $R(\bm{\tau})=0$ for ties.
We design the reward function $r$ with the following heuristic rules:
  • For each episode $\bm{\tau}=({\bm{s}}_{0},{\bm{s}}^{\prime}_{1},{\bm{s}}_{1},\dots,{\bm{s}}^{\prime}_{T},{\bm{s}}_{T})$, the attacker reward and the defender reward at $t$-th turn have $r({\bm{s}}_{t-1},{\bm{u}}_{t})=-r({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})$, so that the game is zero-sum:
    $\sum_{t=1}^{T}r({\bm{s}}_{t-1},{\bm{u}}_{t})+\sum_{t=1}^{T}r({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})=0.$
  • The attacker total reward $R(\bm{\tau})>0$ if the attacker wins, $R(\bm{\tau})<0$ if the defender wins, and $R(\bm{\tau})=0$ if there is a tie.
  • Actions closer to the end of the game should have larger reward weights. Because they are semantically more related to the defender’s guesses or mistakes, which have larger impacts on the final game outcome. Hence, we introduce a decay weight $\gamma\in(0,1)$ such that $\gamma\cdot r({\bm{s}}_{t},{\bm{u}}_{t+1})=r({\bm{s}}_{t-1},{\bm{u}}_{t})$ and $\gamma\cdot r({\bm{s}}^{\prime}_{t+1},{\bm{v}}_{t+1})=r({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})$. Then $\{r({\bm{s}}_{t-1},{\bm{u}}_{t}\}_{t=1}^{T}$ and $\{r({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})\}_{t=1}^{T}$ become two geometrical sequences whose importance enlarges when dialogue turns increase.
  • To improve the LLM training stability, we normalized the total reward $R$ to have norm $|R(\tau)|=1$ if $R(\bm{\tau})\neq 0$.
Based on the above rules, given a complete trajectory, we can assign the reward for each action:
$\left\{\begin{array}[]{lll}r({\bm{s}}_{t-1},{\bm{u}}_{t})=(1-\gamma)\gamma^{T-t}/(1-\gamma^{T+1}),&r({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})=-r({\bm{s}}_{t-1},{\bm{u}}_{t}),&\text{if attacker wins.}\\ r({\bm{s}}_{t-1},{\bm{u}}_{t})=-(1-\gamma)\gamma^{T-t}/(1-\gamma^{T+1}),&r({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})=-r({\bm{s}}_{t-1},{\bm{u}}_{t}),&\text{if defender wins.}\\ r({\bm{s}}_{t-1},{\bm{u}}_{t})=0,&r({\bm{s}}^{\prime}_{t},{\bm{v}}_{t})=0,&\text{if game is tied.}\end{array}\right.$
Note that the above reward design naturally encourages the players to interact less to win the game.
Appendix C Evaluation Benchmark Details
To ensure that our evaluation can be replicated, we use a publicly available code repository called Language Model Evaluation Harness [Gao et al., 2023] to run evaluations on the following benchmarks. We run the evaluation with “dtype=float16” and “max_gen_toks=1024” to make the evaluation faster and more stable. Specifically, we have modified the filtering parameters of BBH by enabling “remove_whitespace” to remove the spaces at the beginning of the model output, allowing the program to better recognize the responses from the model output. We report 5-shot results for MMLU, 3-shot for BBH and 0-shot results for all other benchmarks.
  • MMLU [Hendrycks et al., 2020] is a massive multitask test consisting of multiple-choice questions from various branches of knowledge, requiring the model to possess extensive world knowledge and problem-solving ability.
  • BIG-Bench Hard (BBH) [Suzgun et al., 2022] consists of 23 tasks where humans perform better than language models, which are sourced from BIG-Bench [Srivastava et al., 2022].
  • Mutual [Cui et al., 2020] is a dataset for multi-turn dialogue reasoning modified from Chinese high school English listening comprehension exams.
  • ARC [Clark et al., 2018] question set consists of 7,787 questions and is partitioned into challenge set ARC-c and simple set ARC-e, which require strong knowledge and reasoning abilities to solve.
  • LogiQA2 [Liu et al., 2023] is a revision and re-annotation of LogiQA, a logical reasoning reading comprehension dataset. LogiQA2 has increased data volume, utilized manual translation, and removed cultural features.
  • WinoGrande [Sakaguchi et al., 2019] is a commonsense reasoning dataset consisting of 44k questions with binary options.
  • PIQA [Bisk et al., 2020] is a dataset designed to test the physical commonsense reasoning ability of LLM.
Appendix D Advantange Value Estimation
We estimate the advantage as the TD residual:
$\hat{A}_{\bar{\theta}}({\bm{s}}_{t-1},{\bm{u}}_{t})=r({\bm{s}}_{t-1},{\bm{u}}_{t})+V_{\bar{\theta}}({\bm{s}}_{t})-V_{\bar{\theta}}({\bm{s}}_{t-1})$ (15)
$V_{\bar{\theta}}({\bm{s}})$ is the value function based on the self-play of $\mu_{\bar{\theta}}$ and $\nu_{\bar{\theta}}$, which is based on the language and reasoning ability of the same LLM $\pi_{\bar{\theta}}$. Therefore, we make a reasonable approximation that for all states, the value function of self-plays of the same LLM is equal to 00, which means $\hat{A}_{\bar{\theta}}({\bm{s}}_{t-1},{\bm{u}}_{t})\approx r({\bm{s}}_{t-1},{\bm{u}}_{t}).$
Appendix E Self-play Examples
Target Word momentum
Attacker let’s begin. When you think about physics, what comes to mind?
Defender Gravity, force, energy, and motion are some of the things that come to mind when I think about physics
Attacker Great! Now, let’s consider a scenario where a ball is moving at a constant speed. What would you say is the property of the ball that keeps it moving in the same direction?
Defender The property of the ball that keeps it moving in the same direction is its momentum.
Winner Attacker
Target Word panda
Attacker let’s begin. I’m thinking of a creature that’s native to China and is known for its distinctive black and white coloration. Can you guess what it is?
Defender I’m not sure yet, could you provide more clues?
Attacker Absolutely, I can provide more clues. This creature is a symbol of peace and friendship in many countries. It’s also a popular mascot for various organizations and products. Any ideas?
Defender I’m still not certain. Could you give me another hint?
Attacker Of course, I can give you another hint. This creature is often seen in zoos and is known for its love of bamboo. It’s also a popular choice for children’s books and cartoons. What do you think it might be?
Defender I know the word! It is panda.
Winner Defender
Table 3: Self-play examples of the IM model based on LLaMA-2-7B.
Target Word arbitration
Attacker let’s begin. Imagine a scenario where two parties have a dispute and they don’t want to go to court. What process might they use instead?
Defender They might try to resolve their dispute through mediation or negotiation.
Attacker That’s correct, but what if the dispute is more complex and involves multiple parties? There’s another process that’s often used in such cases. It’s similar to mediation, but it’s more formal and involves a neutral third party who makes a binding decision. Can you name it?
Defender I know the word! It is arbitration.
Winner Defender
Target Word demand
Attacker let’s begin. When you think about the economy, what’s the first thing that comes to your mind?
Defender The first thing that comes to my mind when I think about the economy is employment.
Attacker That’s a good point. Now, considering employment, what would you say drives the need for certain jobs or services in the market?
Defender The demand for goods and services in the market drives the need for certain jobs or services in the economy.
Winner Attacker
Table 4: Self-play examples of SPAG-1 model based on LLaMA-2-7B.
Target Word vaccine
Attacker let’s begin. So, I was reading an article about public health recently. It mentioned how important it is to protect ourselves from certain diseases. What do you think is one of the most effective ways to achieve this?
Defender One of the most effective ways to achieve this is through vaccination.
Attacker That’s correct! Now, can you tell me what we call the substance that is injected into our bodies to help us fight off certain diseases?
Defender I know the word! It is ”vaccine”.
Winner Defender
Target Word credit
Attacker When you go shopping and use a card instead of cash, what do you usually get in return?
Defender I usually get a receipt in return when I use a card instead of cash.
Attacker That’s correct. Now, when you have a good track record of paying back your debts on time, what does that give you access to?
Defender That gives me access to credit cards and loans.
Winner Attacker
Table 5: Self-play examples of SPAG-2 model based on LLaMA-2-7B.
Target Word ray
Attacker When you think about the sun, what comes to mind?
Defender When I think about the sun, I think about warmth, light, and energy.
Attacker That’s interesting. Now, when you think about the sun’s energy, how do you visualize it?
Defender I visualize the sun’s energy as a bright, radiant beam that illuminates everything in its path.
Attacker That’s a great image! Now, when you think about the sun’s energy, what specific part of it do you associate with the light that we see on Earth?
Defender I associate the sun’s energy with the visible light that we see on Earth.
Attacker Great! Now, when you think about the visible light that we see on Earth, what specific part of it do you associate with the light that we see on Earth?
Defender I associate the visible light that we see on Earth with the light that we perceive as white light.
Attacker That’s a good association. Now, when you think about the visible light that we see on Earth, what specific part of it do you associate with the light that we see on Earth?
Defender I know the word! It is ”rays”.
Winner Defender
Target Word eclipse
Attacker What do you think about astronomical events that occur in our solar system?
Defender I find them fascinating. They provide a glimpse into the vastness of space and the intricate workings of our solar system.
Attacker That’s interesting! Can you tell me more about specific celestial events that involve the sun or moon?
Defender Sure, one example is a solar eclipse. This occurs when the moon passes between the sun and Earth, casting a shadow on Earth. Another example is a lunar eclipse, where the Earth passes between the sun and the moon, casting a shadow on the moon. Both events are quite spectacular and can be observed from Earth.
Winner Attacker
Table 6: Self-play examples of SPAG-3 model based on LLaMA-2-7B.
Target Word another
Attacker Can you tell me a synonym for the word ”additional”?
Defender Sure, a synonym for ”additional” could be ëxtra.̈
Attacker That’s correct! Now, can you think of a term that could be used to refer to one more item or person, but isn’t synonymous with ”extra” or ”additional”?
Defender Sure, the term ”supplementary” could be used to refer to one more item or person.
Attacker That’s a good one, but it’s not the term I’m thinking of. This term is often used in everyday conversation and is less formal than ”supplementary”. Can you guess what it might be??
Defender I know the word! It is ”another”.
Winner Defender
Target Word right
Attacker When you’re driving, what do you do when you see a green arrow pointing in a certain direction?
Defender I would proceed in the direction the arrow is pointing.
Attacker That’s correct. Now, imagine you’re reading a book. Which direction do you turn the pages to continue reading?
Defender I would turn the pages to the right to continue reading.
Winner Attacker
Table 7: Self-play examples of GPT-4.